TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes

Create new blog entry for Point 8


Created by Main.MirkoPojer on 2007-05-24 - 19:26

Powering tests in Sector 7-8: activity of Thursday 24/05.

The day was fully devoted to perform tests on Q4, which had experienced an unexpected quench at 300 A on Monday (after ramping at 600 A).

The following tests have been executed on Q4 (MANUAL execution, due to change in procedure):
  • ramp to 600 A with 5 A/s and quench driven by QPS, generating an artificial quench signal from one current lead detector; the test was successful;
  • ramp of B1 power converter to nominal current (3610 A) with 10 A/s ramp rate and simultaneous ramp of B2 power converter to 2000 A with nearly 5 A/s (unbalanced configuration):
    • when B1 current reached 1025 A, its reference current went anormally down to 0;
    • later analysis showed that a fault was sent from the converter and a consequent FPA was started;
    • PO attempt to get control of B2 power converter, to ramp it down contemporarily to 0, provoked a sudden voltage peak in B1, which initiated a quench;
  • at 11.48 (according to PIC supervision) a power cut occured on the general services of UA83, which eventually provoked the loss of controls on cryogenics and lead heating system;
  • after waiting for the good cryo conditions to be recovered, and in agreement with PO, QPS and MPP experts, we decided to ramp again Q4 with the same unbalanced cycle:
    • B1 at 3610 A with 10 A/s ramp rate and B2 to 2000 A with a nearly 5 A/s ramp rate;the test was successful and NO quench was activated;
    • while ramping B2 power converter to the same current as B1 (the idea was to stay on plateau for half an hour and then provoke a quench), and with a current in B2 of 2900 A, a quench generated; preliminary analysis indicated that it was started in B1, possibly due to electromagnetic coupling (this configuration has never been tested before).
  • end of the tests due to long recovery time.

Proposal for the activity of tomorrow.

  • Continue powering Q4 and D2.

  • Continue installation of plastic bags on 600 A current leads.
  • Lower the temperature on the 13 kA leads to check the absence of condensation with AT/MEL copper laminations.

  • ElQA on the selected circuits: list will be provided during the morning.

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-05-24 - 15:02

Powering Tests Sector 7-8 - Minutes 24/05/2007

PDF RAT_20070524b.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x24x14x54/RAT_20070524b.pdf

PowerPoint Q5_test_240507.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x24x14x54/Q5_test_240507.ppt

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-05-24 - 10:06

Sector 7-8 cooldown

After becoming the first magnet at 4.5K and in the kA range to be powered last week, the Q5 matching section quadrupole (MQM) left of Point 8 became the first one to reach nominal current on May 23 @ 15:04:28. The two apertures of the magnet were then powered at different current levels and a quench was provoked. Following that, the magnet was taken to nominal current. This magnet, which was suspected to require training quenches, reached nominal current after two quenches (3910A and 4224 A) before reaching nominal current (4310A) [see the HC Twiki pages for details].

Commissioning now proceeds with Q4 and then D2 as well as the associated orbit correctors the same region.

In the meantime, the cryo team has continued with the tuning of the pumping group and the 4.5K refrigerator and has started cooling down the arc down to 1.9K. This morning (Thursday, May 24) half of the arc was below lambda point. The Power Converter, Interlock, Current Lead, Electrical Quality Assurance and Quench Protection System teams have started the preparation for the powering of the arc early next week.

Comments:

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-05-02 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login